OpenStudy (anonymous):

4096, 1,024, 256, 64, ?,?,?

OpenStudy (anonymous):

Sequence is, we continually divide four to get the same result: \[ A_n=4096\cdot4^{1-n} \]So: \[ 4096\cdot4^{1-5}=\frac{4096}{4^4}=16 \]And: \[ \frac{16}{4}=4 \]For the last one: \[ \frac{4}{4}=1 \]And we are done.